# Rampa: Robotic Augmented Reality for Machine Programming and Automation

### v1

@@ III-A5 Execution on the Real Robot
The execution of the recorded trajectories and the sampled trajectories from the trained model can be done using the virtual user interface. This capability actualizes the transfer of trained tasks from simulation to real-world execution.
III-A6 Collision Detection
Incorporating scene understanding and collision detection, the simulated robot can perceive and interpret its environment. This capability ensures that the robot identifies potential obstacles and issues warnings when a possibility of collusion occurs. This integration of scene understanding, collision detection, and real-world testing ensures the robot operates safely and efficiently, providing confidence in its deployment for various tasks.
@@ III-B Implementation
The implementation can be divided into four parts: In the game engine, where Unity is used, the main functionalities are implemented using Meta Quest API packages. The ROS part manages the interactions between the application and the IK service and the models used for training, along with handling the communication between the application and the actual robot. The actual robot is used to record the demonstrated trajectories and execute the skills effectively upon request. Finally, the XR headset, specifically Quest 3, deploys and operates Unity application where the scene detection features are utilized.
Figure 6: Technical workflow of Rampa. The state of the hand, controller, and room environment is retrieved from the tracking system of the XR headset. The Unity application also runs on the headset.
@@ III-B1 Unity
The simulated robot is integrated into Unity, using the URDF Importer package of Unity, where the simulated robot replicates the kinematics, dynamics, and physical properties according to the Unified Robot Description Format (URDF) file of the robot. The robot also behaves like a physical object and can collide with object meshes obtained from Scene API of Meta Quest packages. A menu is also implemented, enabling the user to draw and adjust trajectories, switch between different trajectory capture modes, and perform the previously mentioned ML tasks.
Detecting and recording the hand movements are done by the use of Meta Quest’s hand gesture recognition and the related APIs. The user can choose whether the positions in the trajectory are recorded with a fixed orientation or if the orientation of the user’s hand is also captured. The user can also choose whether the simulated robot demonstrates the trajectory after the hand-drawn trajectory is completed or if the robot demonstrates it in real time as the trajectory is being drawn.
If the user opts for real-time hand-following by the simulated robot, the IK solver continuously computes the next state of the trajectory, while the ROS client simultaneously sends this state information to Unity for execution as the user continues to draw the trajectory as shown in Fig. 1. Otherwise, the whole data of the trajectory is sent to the ROS client to calculate the necessary joint angles for execution. To facilitate data playback, the number of joint states is kept equal to the number of target points in the trajectory recorded. After necessary joint states are received from the ROS client, the trajectory is ready to be executed with user input. If the simulated robot collides with any object from the surroundings, a warning is triggered to the user. With trajectory data being stored, the user then can playback and adjust the trajectory by redrawing from the current waypoint, using the menu illustrated on the right-hand side of Fig. 4. Trajectory points and orientations can be saved, viewed -which corresponds to yellow trajectories in Fig. 5, and be used to train an ML model in the ROS client upon user request. The trained model can also be conditioned with position and orientation, using virtual cubes, as demonstrated in the right-hand side of Fig. 5.
@@ III-B2 ROS Client
When the Unity client sends a request, including the current state of the robot and target positions, the ROS server responds to the request with a trajectory consisting of joint states, using the solution from the IK solver. Recorded trajectories can be saved and ML models can be trained using these trajectories upon service requests to ROS client.
@@ III-B4 XR Headset
The Unity application is deployed on the headset for complete operation with AR experience. For the co-location of the simulated robot with the actual robot, the left controller’s position and rotation are used with relevant offset values, as illustrated in Fig. 3. For collision detection, Quest 3’s Space Setup feature is utilized, where the current room is scanned with the headset, creating scene anchors, which will then be used in the application as physical meshes.
@@ IV Evaluation
We conducted a comprehensive evaluation using quantitative metrics to assess the competence and performance of the Rampa system and compare it with Kinesthetic Control (KC). These assessments include measuring the time required for task completion, evaluating the performance of trained MP, the smoothness of demonstrated trajectories, and user feedback concerning usability, user experience, and task load through questionnaires.
IV-A Participants
A total of 20 participants (18 male and 2 female, ages 20-36, M=23.5, SD=3.5) were recruited for the user study. Participants include 10 people with experience in robotics, where 7 of them additionally had experience in AR (experts), and 10 people with no prior experience in AR and robotics (novices). Permission for the user study is taken from the Institutional Review Board for Research with Human Subjects of Bogazici University (No:2024/22).
IV-B Experimental Setup
@@
Figure 7: Comparison of Task Completion Times of The Rampa and Kinesthetic Control. The box shows the interquartile range (IQR), including the middle 50% of the data between the first quartile (Q1) and the third quartile (Q3). The whiskers extend from the edges of the box to 1.5 IQR of the Q1 and Q3, and the rest of the data points are considered outliers, shown as red points.
IV-C Procedure
We arranged the user study so that half of the participants first used the Rampa system and then moved the robot in KC mode, while other participants performed these tasks in the opposite order.
Before participants started performing tasks, we offered an overview and explanation of the procedure and experiment. We demonstrated to the user how to use both Rampa system and the physical robot. After participants became familiar with using the Rampa system and moving the real robot in KC mode, we asked them to perform three main tasks:
@@ IV-C Procedure
Participants were requested to draw a trajectory where when the simulated robot mimics the route drawn, it can pass a ball that it holds through a basketball hoop. The robot initially holds a ball and releases it at the end of the route. We also asked them to perform this task by manipulating the real robot with their hands.
@@
Inspired by crane operations for logistics in the industry, a chain hanging from the ceiling and an object on the table with an attached hook were prepared. Users were asked to draw a trajectory for the robot to perform a sequence of actions: moving the chain towards the object, connecting the chain with the hook, lifting the object from the table, and subsequently placing the object back on the table. They were also asked to do the same task by moving the real robot manually with their hands.
@@
We recorded the completion time for each task. In tasks 1 and 2, the duration for demonstrating a single robotic trajectory was recorded. In task 3, we measured the time required to demonstrate three distinct robot trajectories. In task 3, we also saved the demonstrated trajectories for subsequent analysis, allowing for the comparison of their respective smoothness and the similarity among trained MPs. Upon completion of the tasks, the participants were also asked to complete a survey.
IV-D Results & Discussion
@@
In the KC condition, Table I shows that the task completion time difference between expert and novice users was 22.6% for Task 1 and 34.5% for Task 2, indicating a notable difference due to the increased complexity of the demonstration for Task 2 as expected. On the other hand, when utilizing the Rampa system, the task completion time difference was remarkably reduced. Our analysis thus suggests that Rampa allows users to demonstrate robot trajectories efficiently and lessens the impact of prior experience on task performance.
TABLE I: The Comparison of Task Completion Times
KC RAMPA
[…]
@@
To compare the trajectory generation performances of the models learned from Rampa and KC modes, a ground truth trajectory is required for the new environment where these models are tested. For this, we initially trained a contextual ProMP model using the trajectories demonstrated during Task 3, using the heights of the objects as the model’s parameter. We considered the trajectory sampled from this model as the ground-truth trajectory. Then, we computed the Mean Squared Error (MSE) between the ground-truth trajectory and the generated trajectory from the trained and conditioned ProMP by the user during task 3. The result shows that the mean of MSE values is 0.0017 (SD = 0.0024) for the KC condition and 0.0014 (SD = 0.0011) for Rampa condition, and indicates that the ProMP model shows comparable performance when trained on trajectories derived using either Rampa system or kinesthetic control.
@@
We conducted an analysis using the following metrics: jerk, deviation, and variation to analyze the smoothness of trajectories. As shown in Table II, trajectories recorded using the Rampa system show superior or comparable smoothness across the metrics. Our analysis indicates that Rampa facilitates the demonstration of robotic trajectories by increased efficiency while maintaining the consistency and quality of demonstrations.
TABLE II: The Comparison of Smoothness Parameters of Normalized Demonstration Trajectories.
KC RAMPA
[…]
@@
The survey consists of System Usability Scale (SUS) [5], the shorter version of the User Experience Questionnaire (UEQ) [6], the NASA Task Load Index (NASA-TLX) [7], and additional questions to compare the different aspects of Rampa system and KC. We utilized the SUS, Fig. 8, to evaluate the usability of the Rampa system, and we obtained a score of 82.75, which shows that the Rampa system is within A grade (i.e., between 90% and 95%) [31].
Figure 8: SUS Results of The Rampa.
@@
We employed a subset of questions from the SUS to compare the Rampa and KC conditions, and the results are generally favorable towards the Rampa system. As shown in Table III, the Rampa system has superior scores across positive statements such as Q1, Q3, Q5, and Q7, which suggests that the Rampa system provides an easy-to-use framework for demonstrating robotic trajectories. In addition, the reported scores for Q4, Q5, and Q6 suggest that manually moving the robot might be perceived as more demanding and complex compared to using the Rampa system. These findings support that the Rampa enhances the user experience and satisfaction while simplifying the robotic trajectory demonstration process.
TABLE III: SUS Results Comparison of Kinesthetic Control and RAMPA.
Question KC RAMPA
I think that I would like to use
this system frequently.
[…]
I found the system unnecessarily
complex.
[…]
I think that I would need the support
of a technical person to be able to use
this system.
[…]
I would imagine that most people would
learn to use this system very quickly.
[…]
I found the system very cumbersome
to use.
[…]
I needed to learn a lot of things before
I could get going with this system.
[…]
We additionally asked the participants to rate the flexibility of the Rampa system and their learning curve on a scale of 0-5. The results indicate that the average score of flexibility is 4.15 (SD = 0.73), and the average learning curve score is 4.15 (SD = 0.65). High scores imply that users grasped Rampa quickly, and emphasize its user-friendly and intuitive design.
@@
To evaluate the perceived safety of the Rampa system, participants were asked to express which method they found safer. Reported responses were overwhelmingly positive towards the Rampa system, with 95% of participants (19 out of 20) indicating that using Rampa is safer than KC. This result suggests that Rampa addresses safety concerns in human-robot interactions, especially for novice users.
@@
We employed the shorter version of the UEQ to further evaluate the user’s experience of using the Rampa system. As shown in Fig. 9, the Rampa system has positive evaluations across both pragmatic and hedonic qualities. Particularly, Rampa received a score of 2.09 on pragmatic quality and 1.85 on hedonic quality scales, which puts the Rampa system in the range of the top 10% results [6].
Figure 9: UEQ Results of The Rampa.
Figure 10: Comparison of NASA-TLX Results of The Rampa and Kinesthetic Control. The box-and-whisker plot is explained in Fig. 7
@@
The analysis of the NASA-TLX survey with a paired t-test reveals that participants using Rampa were more satisfied with their task performance (t = 2.08, p = 0.044) and felt less frustration (t = -2.518, p = 0.016) while indicating that Rampa requires less physical demand (t = -6.92, p <0.001) and effort (t = -5.04, p <0.001 (Fig. 10). The increased performance satisfaction with reduced physical demand, effort, and frustration suggests that Rampa provides an intuitive and user-friendly interface to perform tasks effectively. These findings show that Rampa enhances not only task efficiency but also user experience and comfort.
@@
At the end of the survey, participants were asked which approach they would prefer to use for demonstrating robotic trajectory. The reported answers show a significant preference towards Rampa with 95% (19 out 20 participants), which, considered with previous evaluation results, indicates that Rampa strongly facilitates the efficacy and accessibility of the robot programming process.
V Conclusion

### v2

@@ III-A5 Execution on the Real Robot
The execution of the recorded trajectories and the sampled trajectories from the trained model can be performed using the virtual user interface. This capability actualizes the transfer of trained tasks from simulation to real-world execution.
@@ III-B Implementation
The implementation can be divided into four parts: In the game engine, where Unity is used, the main functionalities are implemented using Meta Quest API packages. The ROS part manages the interactions between the application and the IK service and the models used for training, along with handling the communication between the application and the actual robot. The actual robot is used to record the demonstrated trajectories and execute the skills effectively upon request. Finally, the XR headset, specifically Quest 3, deploys and operates Unity application. The following sections examine each part in detail, complemented by an illustration of the technical workflow in Fig. 5.
Figure 5: Technical workflow of Rampa.
@@ III-B1 Unity
The simulated robot is integrated into Unity, using the URDF Importer package of Unity, where the simulated robot replicates the kinematics, dynamics, and physical properties according to the Unified Robot Description Format (URDF) description of the robot. A menu is also implemented, enabling the user to draw and adjust trajectories, switch between different trajectory capture modes, and perform the previously mentioned ML tasks.
Detecting the hand movements are done using Meta Quest’s hand gesture recognition and the related APIs. The user can choose from the UI menu provided before starting to demonstrate the trajectory whether the positions in the trajectory are recorded with a fixed orientation or if the orientation of the user’s hand is also captured. The end-effector’s position is determined to be the position of the user hand pinch. If the user chooses to capture orientation, the hand orientation information provided by the related Meta Quest API is used, and the palm-facing-down orientation of the hand is mapped to the orientation in which the end-effector’s tool axis aligns with the negative vertical direction. The user can also choose whether the simulated robot replicates the trajectory after the demonstration or the robot follows the hand in real time.
To enable co-location of the simulated robot with the actual robot, the user places the left controller of the headset to a stable 3D-printed tool. After selecting auto-calibration in the UI, the simulated robot teleports to the position of the controller with a relative offset. For our case, we 3D-printed a tool that can stand on the base of UR10; however, it is important to note that any fixed, stable object that can hold the controller can be used to auto-calibrate, alongside necessary offset calculations.
If the user prefers real-time hand-following with the simulated robot, an off-the-shelf IK solver is continuously queried to compute the trajectory segment from the previous hand location to the current. This is done as the user continues to draw the trajectory as shown in Fig. 2. In this study, the TRAC solver [33] is used for IK computations. Obtained trajectory segments are simultaneously sent to Unity for illustration purposes via the simulated robot. If the user does not want to activate real-time hand-following, then the entire trajectory data, containing uniformly sampled and time-stamped 6D Cartesian positions, is recorded first and then sent to the ROS client at once. In either case, the trajectory sampling rate is fixed at 200 milliseconds. Immediately after receiving waypoints, the ROS client uses the IK solver to compute the necessary joint angles to execute the trajectory upon the user’s request. During execution, a warning is displayed if the simulated robot collides with any object in the surroundings.
Rampa enables users to replay entire trajectories or readjust portions of any length using the menu illustrated on the right-hand side of Fig. 3. It also allows saving trajectories and viewing waypoints, which are shown with yellow lines and black arrows in Fig. 4. One of the most crucial features that Rampa provides is that the recorded data can be used directly to train ML-based models. The user can also condition the ML-based trajectory model by directly observing the actual obstacle and placing a virtual marker accordingly, as shown in Fig. 4. User-specified time-stamps and 6D Cartesian positions of the placed virtual markers are used to condition the model. The output of the conditioned model is then sent to Unity and simulated by the virtual robot for validation purposes.
@@ III-B2 ROS Client
Upon the user’s request, the ROS client reformats the recorded data and trains a new model from scratch. Since neural-network-based frameworks require long training procedures, we demonstrate this feature by training ProMPs.
@@ III-B4 XR Headset
The Unity application is deployed on the headset for complete operation with AR experience. For the co-location of the simulated robot with the actual robot, the left controller’s position and rotation are used with relevant offset values, as illustrated in Fig. 2-A.
@@ IV Evaluation
We conducted a comprehensive evaluation using both quantitative and qualitative metrics to assess the competence and performance of Rampa and compare it with Kinesthetic Control (KC). These assessments include measuring the task-completion time, evaluating the performance of trained MP, the smoothness of trajectories, and user feedback concerning usability, user experience, and task load through questionnaires.
A total of 20 participants (18 male and 2 female, ages 20-36, M=23.5, SD=3.5) were recruited for the user study. Participants include 10 people with experience in robotics, where 7 of them additionally had experience in AR (experts), and 10 people with no prior experience in AR and robotics (novices). Permission for the user study is taken from the Bogazici University (No:2024/22).
@@
IV-A Procedure
We arranged the user study so that half of the participants first used the Rampa system and then moved the robot in KC mode, while other participants performed these tasks in the opposite order. Before participants started performing tasks, we offered an overview and explanation of the procedure and experiment. Specifically, we showed users how they can utilize trajectory playback, edit, save and deletion features of Rampa, if desired or whenever necessary. Regarding KC mode, the users are informed that they will have to demonstrate from scratch, if they desire to modify their demonstrated trajectory. The users are also told that saving and replaying demonstrated trajectory in KC mode will be handled manually by the experimenter. After participants became familiar with using the Rampa system and moving the real robot in KC mode, we asked them to perform three main tasks:
@@ IV-A Procedure
Participants were requested to draw a trajectory where when the simulated robot mimics the route drawn, it can pass a ball it holds through a basketball hoop. The robot initially holds a ball and releases it at the end of the route, which is executed externally by the experimenter.
@@
Inspired by crane operations for logistics in the industry, a chain hanging from the ceiling and an object on the table with an attached hook were prepared. Users were requested to draw a trajectory for the robot to perform a sequence of actions: moving the chain towards the object, connecting it with the hook, lifting the object from the table, and placing the object back on the table.
@@
We recorded the completion time for each task. In tasks 1 and 2, the duration for demonstrating a single successful robotic trajectory and the number of trials needed was recorded. In task 3, we measured the time required to demonstrate three distinct robot trajectories, and also saved the demonstrated trajectories for subsequent analysis, allowing for the comparison of their smoothness and similarity among trained MPs. For training, ProMP is used and the hyperparameter of weights per dimension is set to 20. Upon completion, participants were also asked to complete a survey. During the user study, demonstrations discarded due to technical issues (e.g., headset or robot malfunctions) were not considered.
IV-B Results & Discussion
@@
Figure 6: Comparison of task completion times of Rampa and Kinesthetic Control. The box shows the interquartile range (IQR), including the middle 50% of the data between the first quartile (Q1) and the third quartile (Q3). The whiskers extend from the edges of the box to 1.5 IQR of the Q1 and Q3, and the rest of the data points are considered outliers, shown as red points.
@@
In the KC condition, Table II shows that the task completion time difference between expert and novice users was 22.6% for Task 1 and 34.5% for Task 2, indicating a notable difference due to the increased complexity of the demonstration for Task 2 as expected. On the other hand, when utilizing the Rampa system, the task completion time difference was remarkably reduced. Our analysis thus suggests that Rampa allows users to demonstrate robot trajectories efficiently and lessens the impact of prior experience on task performance.
TABLE II: Comparison of Task Completion Times
RAMPA KC
Task 1 Experts 7.9s $\pm$ 1.3 27.6s $\pm$ 5.2
Novices 7.5s $\pm$ 1.4 33.3s $\pm$ 10.0
Task 2 Experts 17.7s $\pm$ 3.4 69.5s $\pm$ 15.6
Novices 19.2s $\pm$ 4.8 92.1s $\pm$ 18.9
Task 3 Experts 46.8s $\pm$ 7.0 63.4s $\pm$ 5.3
Novices 40.9s $\pm$ 7.0 60.7s $\pm$ 13.0
@@
To compare the trajectory generation performances of the models, a ground truth trajectory is required for the new environment where these models are tested. For this, we initially trained a contextual ProMP model using the trajectories demonstrated during Task 3, using the heights of the objects as the model’s parameter. We considered the trajectory sampled from this model as the ground-truth trajectory. Next, we computed the Mean Squared Error (MSE) between the ground-truth trajectory and the generated trajectory from the trained and conditioned ProMP by the user during task 3. The result shows that the mean of MSE values is 0.0017 (SD = 0.0024) for the KC and 0.0014 (SD = 0.0011) for Rampa, and indicates that the ProMP model shows comparable performance when trained on trajectories derived using either Rampa or KC.
@@
We analyzed the smoothness of trajectories recorded using Rampa and KC. The results reveal that the Rampa demonstrations showed a lower jerk (0.004 ${m}/{s^{3}}$$\pm$ 0.002) compared to the KC demonstrations (0.045 ${m}/{s^{3}}$$\pm$ 0.011) while exhibiting comparable performance in both the deviation (Rampa: 0.111 ${m}$$\pm$ 0.018, KC: 0.109 ${m}$$\pm$ 0.010) and the variation (Rampa: 0.070 ${m^{2}}$$\pm$ 0.021, KC: 0.073 ${m^{2}}$$\pm$ 0.012), implying superior or comparable smoothness across all metrics. This analysis indicates that Rampa facilitates efficient trajectory demonstration while maintaining the consistency and quality of demonstrations.
Number of Trials
All participants successfully completed tasks 1 and 2 in both conditions. For task 1, the number of trials required by each participant was 1, except for one novice user attempting twice while using the Rampa system. For task 2, the average number of trials is 1.111 (SD = 0.314) for expert participants to succeed in both KC and Rampa conditions. For novice users, the mean of the number of trials is 1.2 (SD = 0.4) in KC and 1.4 (SD = 0.49) in Rampa, which is not significantly different (t = -1.5, p = 0.168). A slight increase in the number of trials for task 2 is expected, given the complexity of the demonstration compared to task 1. Our analysis suggests that the number of trials did not show a statistically significant difference between KC and Rampa conditions for both experts and novices in tasks 1 and 2.
@@
The survey consists of System Usability Scale (SUS) [5], the shorter version of the User Experience Questionnaire (UEQ) [6], the NASA Task Load Index (NASA-TLX) [7], and additional questions to compare the different aspects of Rampa and KC. We utilized the SUS to evaluate and compare the usability of Rampa and KC. We obtained a score of 82.75, which shows that the Rampa system is within A grade (i.e., 90%-95%) while KC achieved a score of 62.38 which places KC within D grade [34].
Figure 7: The comparison of SUS results of KC and Rampa conditions.
@@
We employed a subset of questions from the SUS to compare Rampa and KC. As shown in Fig. 7 , the Rampa system has superior scores across positive statements such as Q1, Q3, Q7, and Q9, which suggests that Rampa provides an easy-to-use framework. The reported scores for Q4, Q7, and Q8 suggest that manually moving the robot might be perceived as more demanding and complex compared to using Rampa. These findings support that Rampa enhances the user experience and satisfaction while simplifying the robotic trajectory demonstration process.
@@
To evaluate the perceived safety of the Rampa system, participants were asked to express which method they found safer. Reported responses were overwhelmingly positive towards Rampa, with 95% of participants (19 out of 20) indicating that using Rampa is safer than KC. This result shows increased perceived safety with Rampa.
@@
We employed the shorter version of the UEQ to further evaluate the user’s experience of using the Rampa system. As shown in Fig. 8, the Rampa system has positive evaluations across both pragmatic and hedonic qualities. Particularly, Rampa received a score of 2.09 on pragmatic quality, 1.94 on hedonic quality and 2.01 on overall scales, which puts the Rampa system in the range of the top 10% results [6].
Figure 8: UEQ results of the Rampa condition.
@@
The analysis of the NASA-TLX survey with a paired t-test reveals that participants using Rampa were more satisfied with their task performance (t = 2.08, p = 0.044) and felt less frustration (t = -2.518, p = 0.016) while indicating that Rampa requires less physical demand (t = -6.92, p <0.001) and effort (t = -5.04, p <0.001) in Fig. 9. In addition, we used the RAW-TLX procedure [35] to compute aggregated workload scores. The resulting scores are 27.08 for Rampa and 44.33 for KC, indicating that participants generally perceived a lower workload while using the Rampa system compared to KC condition. The increased performance satisfaction with reduced physical demand, effort, and frustration suggests that Rampa provides a user-friendly interface to perform tasks effectively. These findings show that Rampa enhances not only task efficiency but also user experience and comfort.
@@
Participants were asked which approach they would prefer to use for demonstrating robotic trajectory. The reported answers show a significant preference towards Rampa with 95% (19 out 20 participants), which, considered with previous evaluation results, indicates that Rampa strongly facilitates the efficacy and accessibility of the robot programming process.
Figure 9: Comparison of NASA-TLX results of the Rampa and KC conditions.
Qualitative Feedback
We collected qualitative user feedback from participants to provide a more comprehensive evaluation of Rampa. One participant stated: “The AR system made me feel safe throughout, and I didn’t have to worry about damaging the robot” (P20), highlighting the safety advantages of Rampa. Regarding efficiency, one expert participant noted: “Despite my experience in the domain, it is intimidating even for me to manually control the industrial robots. Rampa simplifies the data collection part of LfD and speeds up my research” (P11). The system’s intuitiveness was highlighted: “The controls with hand gestures being straightforward” (P1) and “AR integration is done seamlessly. It is easy to understand the trajectory of the robot” (P19).
Auto-Calibration Accuracy
To assess the auto-calibration method, we conducted an experiment where we moved the robot to a pre-defined point on the table using the Rampa system. Firstly, we established a reference point by moving the robot manually to that point and recording the end-effector position. Then, we performed auto-calibration, demonstrated a trajectory using the Rampa system, and recorded the end-effector position, which was repeated 10 times. The resulting mean of the position error relative to the reference point is 0.0196 m (SD = 0.0134).
V Limitations
Our system allows collision detection between the simulated robot and the scanned static environment represented with meshes. It currently does not support visual detection of dynamic objects [36], and interaction of the simulated robot with them. For tasks that require intense robot-object interactions, the dynamic objects should be automatically registered as interactable objects to the Unity Engine simulator on-the-fly, which requires realistic modeling of all physical properties such as mass, center of mass, detailed shape and density. While functional, auto-calibration might be time-consuming; easy-to-use solutions like trackable QR codes can facilitate this process [37]. Moreover, controlling the gripper through the AR interface would improve the completeness of the system, diminishing the need for external operations. Also, a visualization of the robot’s workspace can help users grasp its limitations, making the demonstration process less error-prone. Rampa currently advances PbD research by replacing kinesthetic control, lacking a task-centered focus. Future extensions may incorporate egocentric camera views, computer vision, or semantic understanding to mimic daily user interactions and utilize real-world objects. Even though we use the UR10 robotic platform and ProMPs for convenience, we designed Rampa to be robot- and model-agnostic. The system can be adapted to other robotic hardware if their URDF description and its TRAC IK [33] configuration are provided, necessitating only minimal modifications in the Unity application and the ROS client. The system can also be extended by integrating other ML models into the ROS client, if the model can be trained real-time. In addition, although the Unity application depends on Meta Quest APIs, it can be adapted to any headset, given it has a corresponding XR library with hand recognition.
VI Conclusion